If
1-3+4=100
4-6+9=999
2-6+7=879
Then
3-3+6=???

Answer by addingup(3677)

1-3+4=100 so 2=100
4-6+9=999 so 7=999
2-6+7=879 so 3=879
:
We apply this to 3-3+6 (= 6) = (7-3+2 =6) = 999-879+100 = 220 Your answer is 220.